# Large Scale Multi-Task Bayesian Optimization with Large Language Models

### v1

@@ 6 Discussion and Limitations
We first highlight a few limitations. First, our approach not only requires that all tasks in a problem setting have the same input domain (a problem that has been explored e.g. by Fan et al. (2022)). We further require the existence of a task description context $C[f_{t}]$ that can be used in an LLM prompt to define the task. This is likely more difficult, e.g., for hyperparameter optimization, where the primary thing distinguishing tasks is the data that the models are to be trained on; for this setting, approaches such as (Wang et al., 2024b) are likely more appropriate. Finally, we note that the cost of LLM fine-tuning is significantly higher than simple gradient updates of a shared feature extractor.
Despite these limitations, in two real-world applications where BOLT was applicable it yielded strong results. Few-shot generation matched “from scratch” BO runs, and initializing BO from the LLM samples often improved performance further. Moreover, the interplay between the LLM and Bayesian optimization is noteworthy. Despite interest in using LLMs for optimization (Yang et al., 2024; Zelikman et al., 2024; Shypula et al., 2024; Wang et al., 2024a; Madani et al., 2023), finding initial strong solutions to fine-tune them is challenging in some domains. Bayesian optimization, by offering in-depth search, is an excellent candidate for this.
Impact Statement
This work includes an application in designing peptides using AI, specifically for combating antimicrobial-resistant bacteria—a critical challenge in modern medicine. While our work aims to accelerate the development of peptides that can address the growing threat of antimicrobial resistance, we acknowledge the potential for misuse of any “AI for drug design” method to develop harmful biological agents rather than useful drugs (Urbina et al., 2020).
Appendix A Prompt for antimicrobial peptide design.
{minted} [breaklines, fontsize=]latex System: You are a specialized assistant that modifies peptide sequences to enhance antimicrobial activity. Make up to 25User: Seed peptide to be modified Assistant: Modified peptide
@@
Appendix B Training details.
B.1 DKT/FSBO implementation details
@@
For the antimicrobial peptide design task, we train a PPGPR model using the GPyTorch module with a fully connected network with two hidden layers, each with a dimension of 256. We use a batch size of 128, a learning rate of 0.01, and 1024 inducing points for all peptide design experiments.
@@
For the database query plan optimization task, we instead use a fully connected network with two hidden layers, each with a dimension of 64. We use a batch size of 16, a learning rate of 0.01, and 1024 inducing points for all database query plan optimization experiments. Both models were trained for 20 epochs.
B.2 LLM self-augmentation details
@@
For the antimicrobial peptide design task, in each round of self-augmentation, we generate 200 samples for each of the first $800$ train peptides, and add all peptides with a predicted MIC that is lower than $8$ (signaling significant antimicrobial activity) to the training set for the next round of BOLT.
@@
For the database query plan optimization task, we generate 10 samples for each of the 2933 train queries and add all query plans with a running time that is lower than the best plan generated by BAO to the training set for the next round of BOLT.

### v2

@@ 6 Discussion and limitations
We first highlight a few limitations. First, our approach not only requires that all tasks in a problem setting have the same input domain (a problem that has been explored e.g. by Fan et al. [6]). We further require the existence of a task description context $C[f_{t}]$ that can be used in an LLM prompt to define the task. This is likely more difficult, e.g., for hyperparameter optimization, where the primary thing distinguishing tasks is the data that the models are to be trained on; for this setting, approaches such as [46] are likely more appropriate. Finally, we note that the cost of LLM fine-tuning is significantly higher than simple gradient updates of a shared feature extractor.
Despite these limitations, in two real-world applications where BOLT was applicable it yielded strong results. Few-shot generation matched “from scratch” BO runs, and initializing BO from the LLM samples often improved performance further. Moreover, the interplay between the LLM and Bayesian optimization is noteworthy. Despite interest in using LLMs for optimization [49, 51, 36, 45, 21], finding initial strong solutions to fine-tune them is challenging in some domains. Bayesian optimization, by offering in-depth search, is an excellent candidate for this.
Acknowledgments and Disclosure of Funding
N. Maus was supported by the National Science Foundation Graduate Research Fellowship; J. R. Gardner was supported by NSF awards IIS-2145644 and DBI-2400135; C. de la Fuente-Nunez was supported by NIH grant R35GM138201, and by Defense Threat Reduction Agency grants HDTRA11810041, HDTRA1-21-1-0014, and HDTRA1-23-1-0001.
Appendix
Appendix A Prompt details.
Figures Figure 4 and Figure 5 illustrate the prompt templates used for generating optimized query plans and peptide sequences, with GPT-4o-mini-0718. Figure 4 shows the template for database query optimization, where the system acts as an assistant providing efficient join orderings for a given SQL query. Figure 5 displays the template for antimicrobial peptide design, where the system’s role is to modify peptide sequences to enhance antimicrobial activity.
System: You are a helpful assistant that provides efficient join orderings for given queries. User: {SQL query to be optimized} Assistant: {Optimized query plan}
System: You are a specialized assistant that modifies peptide sequences to enhance antimicrobial activity. Make up to 25% sequence modifications based on known antimicrobial peptide properties such as: positive charge, hydrophobicity, and amphipathicity. User: {Seed peptide to be modified} Assistant: {Modified peptide}
@@
Appendix B Compute details.
This section provides details about all computing resources used to run all experiments and create all results provided in this paper. For experiments, we use GPU compute, using a combination of 18 GPUs from our internal compute cluster. The internal cluster has 2 machines with 18 total GPU nodes. One machine has 8 NVIDIA RTX A6000 GPU with 48GB of VRAM, and has 2 sockets with 48 logical threads per socket. The other machine has 10 NVIDIA RTX A5000 GPUs with 24 GB of VRAM, and has 2 sockets with 24 logical threads per socket. Running all experiments needed to produce all results in this work required roughly $25000$ total GPU hours.
Appendix C Implementation details.
C.1 DKT/FSBO implementation details.
@@
For the antimicrobial peptide design task, a PPGPR model was trained using the GPyTorch module. This model employed a fully connected network with two hidden layers, each having a dimension of 256. Training parameters included a batch size of 128, a learning rate of 0.01, and 1024 inducing points for all peptide design experiments.
@@
For the database query plan optimization task, the PPGPR model utilized a fully connected network with two hidden layers, each with a dimension of 64. A batch size of 16, a learning rate of 0.01, and 1024 inducing points were used for these experiments.
For both tasks, 50 STBO optimization trajectories were randomly selected. The DKT/FSBO-$10/20/50$ models were trained using the first 10, 20, or 50 of these trajectories, respectively. All models were trained for 20 epochs.
C.2 POGPE/SGPE implementation details.
Similarly to Section C.1, for the antimicrobial peptide design task, each expert model was a PPGPR model implemented with GPyTorch, with a fully connected network with two hidden layers, each with a dimension of 256. A batch size of 128, a learning rate of 0.01, and 1024 inducing points were used.
For the database query plan optimization task, each expert model uses a fully connected network with two hidden layers, each with a dimension of 64. The training used a batch size of 16, a learning rate of 0.01, and 1024 inducing points.
The same 50 STBO optimization trajectories from Section C.1 were used, and the first $5/10/20$ trajectories were used to train the POGPE/SGPE expert models. In POGPE, all experts were weighted equally. For SGPE experiments, the weighting scheme from Schilling et al. [34] was adopted, where the independent GP for the target dataset carries the same weight as the entire set of experts.
C.3 Optformer implementation details.
For both the query plan optimization and antimicrobial peptide design tasks, GPT-4o-mini-0718 was fine-tuned on past optimization trajectories. To stay within context window limits, a maximum input context length of 100 trials and an output of 20 trials were used. The objective value ranges for both tasks were discretized into 1000 equidistant points. The training sets were constructed by randomly subsampling two trajectories of length 120 from the optimization trajectories. The query plan optimization task is trained on 27.4 million tokens and the antimicrobial peptide design task is trained on 4.8 million tokens. Both models were trained for 1 epoch with a batch size of 20 and an OpenAI learning rate multiplier of 1.8.
Optimization was initialized using the same points as single-task BO. During inference, a constant temperature of 0.7 was used. To manage inference token usage, a batch size of 20 was employed, where the model predicted the next 20 trials based on the previous 100 trials. This was important as experiments ran for 4,000 (query plan) or 20,000 (peptide design) trials.
C.4 LLAMBO implementation details.
The end-to-end LLAMBO method was utilized, leveraging GPT-4o-mini-0718 for several components: generating candidate solutions, serving as a surrogate model for the objective function (via in-context learning), and acting as a conditional sampler to generate candidates for specific target values. Similar to Optformer, a maximum input context window of 100 trials was enforced to prevent exceeding context limits.
The hyperparameters from the original LLAMBO paper were adopted, including an exploration hyperparameter $\alpha=0.1$, and $M=20$. For the surrogate model, we sample $K=10$ MC predictions to compute the empirical estimates. Consistent with the LLAMBO paper, we use the same sampling parameters with a temperature of $0.7$ and top_p of $0.95$. A limit of 10 million maximum input tokens per experiment was used to manage computational costs.
C.5 LLM self-augmentation details.
@@
For the antimicrobial peptide design task, 200 samples were generated for each of the initial $800$ training peptides during each self-augmentation round. Any peptides with a predicted MIC below 8 (indicating significant antimicrobial activity) were added to the training set for the subsequent round of BOLT.
@@
For the database query plan optimization task, 10 samples were generated for each of the 2,933 training queries. Query plans with a runtime lower than the best plan generated by the BAO optimizer were added to the training set for the next round of BOLT.
Appendix D Additional abalations.
D.1 Open source LLMs.
To explore the viability of open-source models for BOLT, QWEN-2.5-7B and LLAMA-3.1-8B were fine-tuned using the identical dataset that created BOLT-1426 from GPT-4o-mini-0718 for the database query optimization task. For evaluation, 50 query plans were generated from each LLM using a sampling temperature of 0.7. The best summed query execution time across the validation tasks from these 50 samples was compared.
Both models were fine-tuned on 4 NVIDIA RTX A6000 GPUs using a per-device batch size of 4, a learning rate of 1e-5 with the AdamW optimizer, and 5 training epochs. The results, shown in Table 3, indicate that QWEN-2.5-7B performed slightly worse than fine-tuned GPT-4o-mini-0718, while LLAMA-3.1-8B showed significantly lower performance. Due to the extensive number of inference calls and multiple fine-tuning rounds required by BOLT, the primary experiments were conducted using the OpenAI API due to hardware resource limitations.
Model Summed runtime
GPT-4o-mini-0718 61.54
QWEN-2.5-7B 62.04
LLAMA-3.1-8B 155.55
Table 3: Comparing open source LLMs fine-tuned with data used to fine-tune BOLT-1426 against OpenAI models fine-tuned on the same data.
D.2 Initializing with previous solutions.
We investigate whether utilizing the best solutions from previously encountered problems provides a sufficiently strong starting point, potentially removing the need for BOLT’s specialized initialization. To assess this, we collect the best-performing solution from each task previously optimized by BOLT-1426. This collection of "previous solutions" is then evaluated on the first 10 validation tasks. Their aggregated performance (summed runtime) is compared against the initialization provided by BOLT (specifically, BOLT-1426, BOLT-1138, and BOLT-893), a standard Bayesian Optimization baseline (STBO), and BOLT-1426 when further refined by a full BO run (BOLT-1426 + BO). The comparison shows that initializing with "previous solutions" yields a higher summed runtime than any of the BOLT initializations or the STBO baseline, suggesting that simply reusing prior optimal solutions is less effective than BOLT’s approach. The results are presented in Table 4.
Method Summed runtime
BOLT-1426 + BO 6.43
BOLT-1426 7.21
BOLT-1138 8.65
BOLT-893 9.08
STBO 8.23
previous solutions 9.18
Table 4: Comparison of summed runtimes on the first 10 validation tasks when initializing with different strategies. "BOLT-1426 + BO" represents full Bayesian optimization initialized by BOLT-1426. "BOLT-1426", "BOLT-1138", and "BOLT-893" show the performance of BOLT’s initialization alone. "STBO" is a standard Bayesian optimization baseline. "Previous solutions" refers to using the best known solutions from all tasks previously seen by BOLT-1426 as the initialization set.
Appendix E Broader Impacts.
This research introduces BOLT, a method leveraging large language models to enhance multi-task Bayesian optimization, with demonstrated applications in antimicrobial peptide design and database query optimization. The potential to accelerate the discovery of novel peptides could significantly benefit public health, particularly in combating antimicrobial resistance. Similarly, improving database query efficiency can lead to substantial computational and energy savings across many industries.
However, we acknowledge the potential for AI misuse in biological design. In applying these powerful methods, expert oversight, rigorous validation, and adherence to established safety and regulatory frameworks must be highlighted. Additionally, the use of large-scale LLMs raises considerations regarding computational accessibility and responsible AI development.